OpenStudy (anonymous):

The parametric equations of a curve are x=2cos t and y=2sin t, for 0 <= t <2 pi. What is the value of t at the point (0,2)?

OpenStudy (anonymous):

\[x=2\cos ~t\]\[y=2\sin~t\] \[0 \le t < \pi\]

sam (.sam.):

You taking A levels?

OpenStudy (anonymous):

Yes. How did you know?

sam (.sam.):

got the feel of it

sam (.sam.):

ok, so just differentiate x and y

sam (.sam.):

x=2cost dx/dt=-2sint -------------------- y=2sint dy/dt=2cost

sam (.sam.):

then \[\frac{dy}{dx}=\frac{2cost}{-2sint}\] \[\frac{dy}{dx}=-\cot~t\]

sam (.sam.):

\[y-y_1=m(x-x_1)\] \[y-2=-\cot~t(x-0)\] \[y=-xcot ~t +2\]

sam (.sam.):

Try Substitute (0,2) into x=2cos t and y=2sin t

sam (.sam.):

t should be pi/2

OpenStudy (anonymous):

Ok, that's the answer. but why?

OpenStudy (anonymous):

Oh, I know why. Thank you! :D

sam (.sam.):

Because x=2cos t and y=2sin t is a part of the curve, and y=−xcot t+2 also part of a curve, So, substitute (0,2) into both equation will get the same value